This is a Mixe worth 1 point / You have 2 of 6 attempts remaining / There is a 396 attempt penalty 12 Question (1 point) See page 569 Hydrogen gas reduces NO to N2 in the following reaction: 2I12(g) +2NO(g211,0g)+N2g The initial reaction rates of four mixtures of H2 and NO were measured at 1500°C with the following results: Experiment | [H2]0 (M) 0.212 0.212 0.424 0.848 | [NOlo (M) 0.136 0.272 0.544 0.544 | Initial Rate (M/s) 0.0303 0.0609 0.480 1.93 2 4 v 5th attempt Part 1 (0.5 point) See Periodic Table < 12/14 VIEW SOLUTION 11OF 14QUESTIONS COMPLETED SUBMIT ANSWER

I need help with part b which is

Determine the rate constant for the reaction at 1500°C.

The answer should be ___________ M^-x S^-1
Answer #1

ustoro K is too rata Constant and χ,y autka ondons δ(stka reachon. hon expuiment 2, o.OF09 _ κ x (o.212)χ x (o . 272)a fem oxpewnout 4; y, 193 = κ x (o. 348)χ k@.54 4)T (g) peu, ment l, KX (D.212) x(0.136)y 0.0303 = () . 12 Sinu Dr, 42 2 do; the Tate ex PTeSSiom is

Similarly from experiment 2, we get :

k = 0.0609/ (0.212)2 * (0.272) M-2s-1 = 4.98 M-2s-1

from experiment 3, we get :k = 0.480/ (0.424)2 * (0.544) M-2s-1 = 4.91 M-2s-1

from experiment 4, we get :k = 1.93 / (0.848)2 * (0.544) M-2s-1 = 4.93 M-2s-1

Average value of k = (4.96 + 4.98+ 4.91+4.93)/4 M-2s-1 = 4.95 M-2s-1